OpenStudy (taepoohwae):

Can someone solve this with the steps? 4x-y=4 9x-2y=11 Use the method of substitution to solve the system of this linear equation. What is the ordered pair?

hartnn (hartnn):

4x-y=4 can you isolate 'y' from this equation ?

OpenStudy (taepoohwae):

I have no idea. My directions are to use the method of substitution to solve the system of this linear equation.

hartnn (hartnn):

Hint: add y on both sides, and then subtract 4 from both sides.

OpenStudy (anonymous):

4x - y = 4 --> 4x - 4 = y now sub 4x - 4 in for y in the 2nd equation. 9x - 2y = 11 9x - 2(4x - 4) = 11 distribute the -2 through the parenthesis and solve for x. Once you have x, then you can substitute that back into the 1st equation and solve for y. You can then sub your known variables in either of the original equations to check your answers. If it comes out equal, then your answers are correct.